### v1

# Quantum Kerr Black Hole from Matrix Theory of Quantum Gravity

Abstract
Recently, a quantum mechanical theory of quantum spaces described by a large $N$ non-commutative coordinates is proposed as a theory for quantum gravity[1]. In this paper, we construct Kerr black hole as a rotating noncommutative geometry solution of this theory. Due to rotation, the fuzzy sphere is deformed into a fuzzy ellipsoid, which matches exactly the outer horizon of the Kerr black hole in the Boyer-Lindquist coordinates. Together with a half-filled Fermi sea, the fuzzy solution reproduces the Bekenstein-Hawking entropy as well as the mass of the Kerr black hole. These results provide further support that the proposed theory of quantum spaces is a plausible candidate for the theory of quantum gravity.
1 Introduction
General relativity is a classical dynamical theory of continuum spacetime. Classical spacetime is described by coordinate functions $x^{\mu}(P)$ defined over the spacetime manifold ${\cal M}$ and $P\in{\cal M}$ is any point (event) in the spacetime. Einstein theory of spacetime consists of an endorsement of a metric on ${\cal M}$. One may allow the use of any coordinate systems and impose diffeomorphism invariance as a symmetry. The simplest Lagrangian which is consistent with the diffeomorphism symmetry is the Hilbert-Einstein action. One then show that this dynamical theory of spacetime contains gravity by finding solutions of the theory and by examining their effects on motion of probes described by geodesic motion.
A particular interesting and puzzling solution is the black hole. First it is not a consistent solution as it contains singularity [2], meaning that the description of black holes in terms of a classical spacetime is incomplete. Besides, it exhibits a Bekenstein-Hawking entropy [3, 4] whose microscopic origin is unknown. (see however [5] for success in microstates counting for supersymmetric black holes in string theory). Nevertheless, the area dependent nature of the entropy suggests that quantum gravity is holographic [6, 7]. Moreover, native consideration of black hole evolution suggests a loss of quantum information [8, 9, 10]. It is believed that only in a theory of quantum gravity can these problems be properly addressed and resolved.
In a recent paper [1], a formulation of quantum gravity in terms of a quantum mechanics of quantum space was proposed. The proposal was motivated by a number of considerations. Recent studies of black hole in the AdS/CFT correspondence [11] has cumulated at the success in obtaining the Page curve behavior [12, 13] of the entanglement entropy of the Hawking radiation. However, the origin and nature of island, a new saddle point in the holographic entropy formula in the island proposal [14, 15, 16, 17], is not clear from a fundamental point of view. Motivated by the curious feature that island is located just underneath the black hole horizon, we postulated [18] that the quantum states of black hole responsible for the Bekenstein-Hawking entropy are given by a thin shell of Bell degrees of fermionic freedom located at the region just underneath the horizon. This picture was further developed in [19, 20] into a bottom-up model of quantum black hole where the energy and entropy of black holes were successfully accounted for.
To make progress, one wants to have a dynamical theory of quantum space where a quantum black hole is obtained as a solution. Our phenomenological model [19, 20] suggests the desired theory must contains fermionic degrees of freedom. Fermions are featured in various popular approaches to quantum gravity such as string/M theory [21], AdS/CFT [11], BFSS quantum mechanics [22, 23] or IKKT matrix model [24]. Since supersymmetry is always built in, the fermion content as well as the supersymmetric action are highly constrained. We argued in [1] that there is actually no essential reason to insist on having supersymmetry if one is interested in quantum mechanics. Gathering the ideas discussed above, we proposed in [1] a $SU(N)$ quantum mechanics of noncommutative geometry as a theory of quantum gravity. The theory has a Higgs-like bosonic potential which allows for the construction of nontrivial “Higgs-vacuum”. This is possible since we do not assume supersymmetry. In general, solutions of the theory are given by dynamical noncommutative geometry. In [1], an interesting noncommutative geometric solution of fuzzy sphere was found. Together with a half-filled Fermi sea residing over the fuzzy sphere, the solution possesses energy of a Schwarzschild black hole and a microstates entropy which agrees precisely with the Bekenstein-Hawking entropy expected for a quantum black hole. It was also shown that the interaction energy between two fuzzy spheres has the correct dependence on the Newton constants and masses as required by gravity. It was conjectured that Newton gravity between static sources would be reproduced in the large distance limit.
In this paper, we construct a new time dependent solution of the quantum mechanics. The solution describes a rotating fuzzy ellipsoid. With a half-filled Fermi sea, the solution has a microstate counting that reproduces precisely the Bekenstein-Hawking entropy of a quantum Kerr black hole. Moreover, the energy matches that of the classical mass of Kerr black hole, with the leading correction at the order of $a^{4}/r_{+}^{4}$. We propose that the fuzzy ellipsoid with a half-filled Fermi sea as a quantum description of the Kerr black hole in quantum gravity. Since a rotating black hole is a much more complicated system, the fact that our proposed dynamical theory of quantum space contains also the Kerr black hole provides further support that the theory is a plausible candidate for the theory of quantum gravity.
2 Kerr Metric in General Relativity
In general relativity, a black hole with mass $M$ and angular momentum $J$ is described by the Kerr metric. The Kerr metric can be written in a number of different ways. See for example [25] for a brief introduction to the Kerr spacetime. In the Boyer-Lindquist coordinates, the Kerr metric reads
$ds^{2}=-\left(1-\frac{2M}{\rho^{2}}\right)dt^{2}-\frac{4Mar\sin^{2}\theta}{\rho^{2}}dtd\phi+\frac{\Sigma}{\rho^{2}}\sin^{2}\theta d\phi^{2}+\frac{\rho^{2}}{\Delta}dr^{2}+\rho^{2}d\theta^{2},$ (1)
where
$\rho^{2}:=r^{2}+a^{2}\cos^{2}\theta,\quad\Delta:=r^{2}-2Mr+a^{2},\quad\Sigma:=(r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta.$ (2)
The Boyer-Lindquist coordinates is a generalization of the Schwarzschild coordinates which allows the use of spherical coordinates. In the asymptotic region $r\to\infty$, we have
$ds^{2}=-\left[1-\frac{2M}{r}+O\left(\frac{1}{r^{3}}\right)\right]dt^{2}-\left[\frac{4Ma\sin^{2}\theta}{r}+O\left(\frac{1}{r^{3}}\right)\right]d\phi dt{}$
$+\left[1-\frac{2M}{r}+O\left(\frac{1}{r^{2}}\right)\right]\left[dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right].$ (3)
This confirms that $M$ is the mass and $J=aM$ is the angular momentum of the black hole spacetime. It is interesting to note that for $M=0$, the metric (1) is flat and reads
$ds^{2}=-dt^{2}+\frac{r^{2}+a^{2}\cos^{2}\theta}{r^{2}+a^{2}}dr^{2}+(r^{2}+a^{2}\cos^{2}\theta)d\theta^{2}+(r^{2}+a^{2})\sin^{2}\theta d\phi^{2}.$ (4)
This is in fact the standard Minkowski space written using the “oblate spherical” coordinates, which is related to the Cartesian coordinates with the coordinate transformation
$x=\sqrt{r^{2}+a^{2}}\sin\theta\cos\phi,\quad y=\sqrt{r^{2}+a^{2}}\sin\theta\sin\phi,\quad z=r\cos\theta.$ (5)
It is clear that, for $a\neq 0$, the oblate spherical coordinates $(r,\theta,\phi)$ is different from the standard polar coordinates for 3-dimensional Euclidean space.
The Kerr metric is a stationary, axially symmetric and asymptotically flat vacuum solution to the Einstein equation. It has the Killing vectors $t^{\alpha}=\partial x^{\alpha}/\partial t$ and $\phi^{\alpha}=\partial x^{\alpha}/\partial\phi$. The metric is invariant under simultaneous inversion of $t\to-t$ and $\phi\to-\phi$, which is consistent with the fact that time reversal of a rotating object reverse the rotation of the object. The metric reveals a curvature singularity $R^{\alpha\beta\gamma\delta}R_{\alpha\beta\gamma\delta}\sim 1/\rho^{12}$ at $\rho=0$. In addition the metric has a coordinate singularity at $\Delta=0$. For $a\leq M$, the roots of $\Delta=0$ are real. The larger solution
$r_{+}=M+\sqrt{M^{2}-a^{2}}$ (6)
denotes the position of the outer horizon. That the surfaces $r=r_{\pm}$ are event horizon can be seen from the fact that they are null: $g^{\alpha\beta}\partial_{\alpha}r\partial_{\beta}r=g^{rr}=\Delta/\rho^{2}=0$ at $r=r_{\pm}$. Examination of the null geodesics near the surface shows that they are indeed horizons. Viewed in the Cartesian coordinates (5), the horizon is an elliptical surface
$\frac{x^{2}+y^{2}}{r_{+}^{2}+a^{2}}+\frac{z^{2}}{r_{+}^{2}}=1.$ (7)
We also remark that the outer horizon is a Killing horizon since the time like Killing vector $\xi^{\alpha}:=t^{\alpha}+\omega_{H}\phi^{\alpha}$ becomes null at $r=r_{+}$, where
$\omega_{H}=\frac{a}{r_{+}^{2}+a^{2}}$ (8)
is the angular velocity of the black hole.
For the Kerr metric, the area of the horizon is given by
$A=4\pi(r_{+}^{2}+a^{2}),$ (9)
which give rises to the Bekenstein-Hawking entropy
$S=\frac{A}{4G}=\frac{\pi(r_{+}^{2}+a^{2})}{G}.$ (10)
The mass of the Kerr black hole, written in terms of $r_{+}$, is
$M=\frac{r_{+}^{2}+a^{2}}{2Gr_{+}}.$ (11)
We will show below that these properties (10) and (11) of the Kerr black hole are reproduced by a rotating solution in our proposed quantum mechanics of quantum space.
3 Kerr Black Hole from Quantum Mechanics of Quantum Space
Consider the proposal of [1] for a $SU(N)$ quantum mechanical theory of 3-dimensional quantized space defined by the Lagrangian
$L=\mbox{tr}\left[\frac{1}{2M_{P}}\dot{X}^{a2}+\frac{M_{P}}{N^{2}}[X^{a},X^{b}]^{2}+\frac{4M_{P}}{N^{2}}X^{a2}+i\dot{\psi}^{\dagger}\psi+\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi-M_{P}{\bf 1}\right],$ (12)
or the Hamiltonian
$H=\mbox{tr}\left[\frac{M_{P}}{2}P^{a2}-\frac{M_{P}}{N^{2}}[X^{a},X^{b}]^{2}-\frac{4M_{P}}{N^{2}}X^{a2}-\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi+M_{P}{\bf 1}\right].$ (13)
Here $X^{a}_{mn}$, $a=1,2,3$ represent the coordinates of a 3-dimensional quantized space, $\sigma^{a}$ are the Pauli matrices and $\psi_{mn},\psi_{mn}^{\dagger}$ are two fermionic coordinates. The theory has $SU(N)$ symmetry as well as a flavor $SO(3)$ symmetry. We do not assume supersymmetry, and so the choice of variables as well as the form of the Lagrangian is not dedicated by symmetry. Instead, we took a bottom up approach in the spirit and art of phenomenological model building. Note that the normalization of the mass scale for the $\dot{X}^{2}$ term was left unfixed in [1] since it is not sensitive to the time independent configurations considered there. Here we adopt a natural choice of scales in (13) where there only appears the Planck mass $M_{P}$. The Planck mass and the Planck length are defined in terms of the Newtonian constant $G$ as
$M_{P}^{2}=\frac{2}{\pi G},\quad l_{P}=\frac{2}{\pi M_{P}}=\sqrt{\frac{2G}{\pi}}.$ (14)
We remark that while the coefficient of the Yukawa term does not affect the results in [1], this becomes important for our analysis here since the fermionic perturbation over the fuzzy ellipsoid is non-vanishing. We show below that the Yukawa term in (13), which together with the bosonic contribution reproduces precisely the energy of the classical Kerr black hole.
Next let us consider solution to the theory. The classical equation of motion for a bosonic matrix configuration is given by
$-\ddot{X}^{a}+\frac{4M_{P}^{2}}{N^{2}}\left([X^{b},[X^{a},X^{b}]]+2X^{a}\right)=0.$ (15)
The Kerr black hole is a stationary solution of general relativity with axial symmetry. Suppose the rotation is around the $z$-axis, then it is natural to consider a new basis of coordinates $(X^{+},X^{-},X^{3})$ where
$X^{\pm}:=\frac{1}{\sqrt{2}}(X^{1}\pm iX^{2}).$ (16)
In the new basis, the equations of motion reads
$-N^{2}\ddot{X}^{+}+4M_{P}^{2}\left([X^{+},[X^{+},X^{-}]]+[X^{3},[X^{+},X^{3}]]+2X^{+}\right)=0,$ (17)
$-N^{2}\ddot{X}^{3}+4M_{P}^{2}\left([X^{-},[X^{3},X^{+}]]+[X^{+},[X^{3},X^{-}]]+2X^{3}\right)=0.$ (18)
The equation of motion for $X^{-}$ is not independent and can be obtained from the one (17) of $X^{+}$. For rotation, let us consider the ansatz
$X^{\pm}(t)=e^{\pm i\omega t}X^{\pm}(0),\quad\mbox{$X^{3}$ independent of $t$}$ (19)
which is appropriate for axial symmetric rotational motion. The equation of motion becomes
$[X^{+},[X^{+},X^{-}]]+[X^{3},[X^{+},X^{3}]]+2c^{2}X^{+}=0,$ (20)
$[X^{-},[X^{3},X^{+}]]+[X^{+},[X^{3},X^{-}]]+2X^{3}=0$ (21)
where $c^{2}:=1+N^{2}\omega^{2}/(8M_{P}^{2})$. The equations (20) and (21) can be solved by utilizing the spin $j=(N-1)/2$ representation of $SU(2)$ in the Cartan-Weyl basis. They satisfy the commutation relations
$[T^{+},T^{-}]=T^{3},\quad[T^{3},T^{\pm}]=\pm T^{\pm}$ (22)
and the Casimir relation
$T^{+}T^{-}+T^{-}T^{+}+T_{3}^{2}=\frac{N^{2}-1}{4}{\bf 1}.$ (23)
It is easy to see that
$X^{\pm}=e^{\pm i\omega t}T^{\pm},\quad X^{3}=c_{3}T^{3},\quad c_{3}=\sqrt{1+\frac{N^{2}\omega^{2}}{4M_{P}^{2}}},$ (24)
solve the equation of motion. As $c_{3}\geq 1$, we can introduce the definition
$\cos\beta=1/c_{3},$ (25)
which is a measures of the angular velocity. Let us introduce the dimensional coordinates $Y^{1,2}=2l_{P}X^{1,2}$, $Y^{3}=2l_{P}X^{3}/c_{3}^{2}$, which are equivalent to
$Y^{1,2}=l_{P}\sqrt{N^{2}-1}\;\hat{T}^{1,2},\quad Y^{3}=l_{P}\sqrt{N^{2}-1}\cos\beta\;\hat{T}^{3}$ (26)
where $\hat{T}^{a}:=\frac{2}{\sqrt{N^{2}-1}}T^{a}$. The reason for this choice of the Cartesian coordinates $Y^{a}$ is such that it matches precisely with the expression (5) of the Cartesian coordinates, which are written in terms of the oblate spherical coordinates. In fact in (26), $\hat{T}^{a}$ can be identified with the directional cosine in (5), and the prefactors can be identified with $R$ and $a$ in (5) if
$a^{2}+R^{2}=l_{P}^{2}(N^{2}-1),\quad R^{2}=l_{P}^{2}(N^{2}-1)\cos^{2}\beta.$ (27)
In terms of $Y$’s, our solution takes the form
$\frac{Y_{1}^{2}+Y_{2}^{2}}{R^{2}+a^{2}}+\frac{Y_{3}^{2}}{R^{2}}=1,$ (28)
together with the commutation relations
$[Y^{+},Y^{-}]=\frac{2R/\cos\beta}{\sqrt{N^{2}-1}}Y^{3},\quad[Y^{3},Y^{\pm}]=\pm\frac{2R\cos\beta}{\sqrt{N^{2}-1}}Y^{\pm}.$ (29)
Note that the relations (28) and (29) are time independent and they define an oblate fuzzy ellipsoid for $\omega\neq 0$. Note also that (27) takes on the simple form
$R=Nl_{P}\cos\beta,\quad a=Nl_{P}\sin\beta$ (30)
in the leading order of large $N$ limit, which we will assume from now on.
Over the geometry of the fuzzy ellipsoid, the Hamiltonian of the system reads
$H=H_{B}+H_{F},$ (31)
where $H_{B}$ is the bosonic part of the Hamiltonian and $H_{F}$ is the part of the Hamiltonian which contains the fermions. It is
$H_{B}=\frac{NM_{P}}{2}+\frac{N^{3}\omega^{2}}{12M_{P}},$ (32)
where we have used $\mbox{tr}T_{3}^{2}=\frac{1}{3}\mbox{tr}T^{2}=\frac{N^{3}}{12}$. The fermionic Hamiltonian $H_{F}=-\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi$ is more interesting. Due to the anisotropy introduced by the coordinates (24), we can write
$H_{F}=H_{F}^{0}+h_{F},$ (33)
as an isotropic part
$H_{F}^{0}=-\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}T^{a}(t)\psi$ (34)
plus a term $h_{F}$
$h_{F}=-(c_{3}-1)\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{3}T^{3}\psi:=\psi^{\dagger}\tilde{h}_{F}\psi,$ (35)
where in (34),
$T^{\pm}(t)=e^{\pm i\omega t}T^{\pm},\quad T^{3}(t)=T^{3}$ (36)
denotes the generators of $SU(2)$ under a rotation of angular velocity $\omega$ around the $z$-axis, and for simplicity, we have introduced the matrix
$\tilde{h}_{F}:=-(c_{3}-1)\frac{M_{P}}{N^{2}}\sigma^{3}T^{3}.$ (37)
The term $h_{F}$ can be considered a perturbation and is of the order of $a^{2}/R^{2}$ when rotation is slow. The unperturbed Hamiltonian $H_{F}^{0}$ can be diagonalized as in [1] and give rises to a Fermi sea. In fact, the matrix $K:=\sigma^{a}T^{a}(t)$ admits the eigendecomposition
$K_{(m\alpha)(n\beta)}=\frac{N}{2}\sum_{p=1}^{N}\left({\cal U}^{p}_{m\alpha}{\cal U}^{p\dagger}_{n\beta}-{\cal V}^{p}_{m\alpha}{\cal V}^{p\dagger}_{n\beta}\right),$ (38)
where ${\cal U}^{p}_{n\beta},{\cal V}^{p}_{n\beta}\;(p=1,\cdots,N)$ are eigenvectors of $K$, which now become time-dependent. They satisfy the orthonormality condition
${\cal U}^{p\dagger}_{m\alpha}{\cal U}^{q}_{m\alpha}={\cal V}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\alpha}=\delta^{pq},\qquad{\cal U}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\alpha}=0$ (39)
and the completeness relation
${\cal U}^{p}_{m\alpha}{\cal U}^{p\dagger}_{n\beta}+{\cal V}^{p}_{m\alpha}{\cal V}^{p\dagger}_{n\beta}=\delta_{mn}\delta_{\alpha\beta},$ (40)
which are however time independent and identical to before. Introducing the fermionic oscillators
$\xi^{p}_{k}:={\cal U}^{p\dagger}_{n\beta}\psi_{nk\beta},\quad\chi^{p\dagger}_{k}:={\cal V}^{p\dagger}_{n\beta}\psi_{nk\beta},$ (41)
$H_{F}^{0}$ can be written in the diagonal form
$H_{F}^{0}=-\frac{M_{P}}{2N}\left(\sum_{p,k=1}^{N}\xi_{k}^{p\dagger}\xi_{k}^{p}+\chi_{k}^{p\dagger}\chi_{k}^{p}-N^{2}\right)$ (42)
and has the eigenstates
$\left|\Psi^{p_{1}\cdots p_{r}q_{1}\cdots q_{s}}_{k_{1}\cdots k_{r}l_{1}\cdots l_{s}}\right\rangle:=\xi^{p_{1}\dagger}_{k_{1}}\cdots\xi^{p_{r}\dagger}_{k_{r}}\chi^{q_{1}\dagger}_{l_{1}}\cdots\chi^{q_{s}\dagger}_{l_{s}}|0\rangle,$ (43)
where $|0\rangle$ is the fermionic Fock vacuum and the eigenvalues
$H_{F}^{0}=\frac{M_{P}}{2N}n,\quad n:=N^{2}-r-s,$ (44)
where $n=-N^{2},\cdots,0,\cdots,N^{2}$ specifies the energy level within the Fermi sea. The lowest level of energy is $-NM_{P}/2$ which corresponds to a completely filled Fermi sea, while the highest energy $NM_{P}/2$ corresponds to an empty Fermi sea.
We can divided the space of eigenstates of $H_{F}^{0}$ according to their corresponding energy level. In general, the set of eigenstates of level $n$ is given by
$\mathbb{V}_{n}:=\{|i\rangle,i=1,\cdots,\Omega_{n}\},$ (45)
where $|i\rangle$ are those states (43) with $r+s=N^{2}-n$. Here
$\Omega_{n}=\binom{2N^{2}}{N^{2}-n}\approx 2^{2N^{2}-|n|}$ (46)
is the degeneracy of the energy level $n$. For a half-filled ($n=0$) Fermi sea, we have
$\Omega_{0}=2^{2N^{2}}$ (47)
in the leading order of large $N$. These microstates give rise to the entropy $S=\log_{2}\Omega_{0}$:
$S=2N^{2}=\frac{\pi(R^{2}+a^{2})}{G},$ (48)
which is precisely the Bekenstein-Hawking entropy of a Kerr black hole if we identify $R$ with $r_{+}$ of the outer horizon radius of the Kerr black hole. To make this identification, we need to show that $R$ actually satisfies the relation (11) for a given mass $M$ of the Kerr black hole. For this, we need to include the energy of the system of Fermi oscillators in the fuzzy background.
For the half-filled Fermi sea, we have $H_{F}^{0}=0$ and we are left with the correction to the energy due to the perturbation $h_{F}$. When the perturbation $h_{F}$ is turned on, we need to use the perturbation theory for degenerate energy level to compute the correction to the $n=0$ energy. This requires the knowledge of the matrix elements $\langle i|h_{F}|j\rangle$ and it’s eigenvalues $\lambda_{p}$, where $|i\rangle,i=1,\cdots,\Omega_{0}:=2^{2N^{2}}$ are the unperturbed states (43) with $r+s=N^{2}$. Rather than computing the individual corrections, it is more meaningful to compute the ensemble average of the corrections
$\overline{h_{F}}:=\frac{1}{\Omega_{0}}\sum_{p}\lambda_{p}$ (49)
of the set of microstates. It was shown in the appendix of [1] that
$\overline{h_{F}}=\mbox{tr}(Kh_{F}).$ (50)
Only the $T^{3}$ component of $K$ contributes and so we get
$\overline{h_{F}}=-\frac{c_{3}-1}{6}NM_{P}.$ (51)
Adding it to (32) and expressing it in terms of macroscopic quantities $R$ and $a$, we have
$E=\frac{R^{2}+a^{2}}{2GR}f(\beta),\quad f(\beta):=\frac{2\cos\beta}{3}\left(1+\frac{1}{\cos^{2}\beta}-\frac{1}{2\cos\beta}\right).$ (52)
For small $\beta$, $f$ has the expansion
$f=1+\frac{1}{6}\sin^{4}\beta+\cdots,$ (53)
which is actually an expansion for small angular momentum since
$\sin\beta=\frac{a}{\sqrt{R^{2}+a^{2}}}.$ (54)
The energy (52) agrees precisely with that (11) of general relativity, with a correction at the order of small $a^{4}/R^{4}$. It is quite remarkable that the Yukawa term in the Lagrangian (12) not only give rises to the needed microstates to account for the entropy of quantum black hole, but also makes contribution to the energy of the fuzzy system, without which the mass of the Kerr black hole would not be obtained. It is interesting to exact spin dependent force between black holes from the quantum mechanics. It is also interesting to think about how the ergosphere of Kerr black hole appears in our description. A probe analysis may help to reveal it. We leave these interesting problems for later consideration.
Acknowledgments
The support of this work by NCTS and the grant 110-2112-M-007-015-MY3 of the National Science and Technology Council of Taiwan is gratefully acknowledged.